Circular No. 7092

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE 1999F, 1999G, 1999H, 1999I A. Kim, Collège de France, Paris, on behalf of the EROS collaboration,

reports the discovery of four apparent supernovae discovered in the course of their automated supernova search conducted with the 1-m Marly telescope at the European Southern Observatory, La Silla. Finding charts can

be found at http://www.lal.in2p3.fr/EROS/SNe/Charts/. All reference

images have limiting magnitudes near mag 22 and were also taken at the Marly telescope.

SN  1998 UT  α_{2000}  δ_{2000}  V  Offset
1999F Jan. 12 11^{h}27^{m}15.^{s}05 − 6^{o}10′33′′.4 19.8: 6′′.6 W, 3′′.3 S
1999G Jan. 12 11 15 39.13 − 5 29 24.8 21.4: 2′′.4 E, 2′′.2 S
1999H Jan. 14 10 07 03.06 − 5 55 26.6 20.6: 1′′.0 E, 0′′.8 N
1999I Jan. 15 10 49 58.09 − 6 14 36.8 21.5: 2′′.2 E, 2′′.7 N

Spectroscopic observations on Jan. 17 with the New Technology Telescope (+ EMMI) show that SN 1999F is a type-Ia supernova, ∼ 1 week after maximum; the host galaxy redshift is 0.18. Spectroscopic observations on Jan. 17 with the NTT show that SN 1999G is a type-Ia supernova at maximum; the host galaxy redshift is 0.17. SNe 1999F and 1999G were not seen on frames taken on 1997 Mar. 30. Photometric confirmations of SNe 1999H and 1999I were made on Jan. 18 at the Marly telescope. SN 1999H was not seen in an image taken on 1997 Mar. 29. SN 1999I was not seen in an image taken on 1997 Apr. 30.

SUPERNOVA 1999C IN ANONYMOUS GALAXY A. V. Filippenko, D. C. Leonard, and A. G. Riess, University of Cal-

ifornia at Berkeley, report that a spectrum (range 510–880 nm) obtained on Jan. 20 UT with the Keck-II telescope reveals that SN 1999C (cf. IAUC 7088) is of type Ia, about two weeks past maximum brightness. Its redshift is 0.125, which differs from that of Abell 914 (0.195).

SUPERNOVA 1998eu IN ANONYMOUS GALAXY

Corrigendum. On IAUC 7082, line 7, for (z = 1.88). read (z = 0.188).

1999 January 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green